# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v11

@@ 2.2 Distributional Zero and Distributional Pole
[…]
$\rho(\tilde{z})d\tilde{z}$= $\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y}$; (2.2)
@@
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, sine $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [5]. A linear transformation generates neither a distributional zero nor a distributional pole, in accordance with to Formula (2.2).
  • Figure 1 illustrates the probability density function for $(x\pm 1)^{2}$ according to Formula (2.3), which exhibits a distributional pole at $x=0$. The distribution $(0\pm 1)^{2}$ corresponds to the $\chi^{2}$ distribution [2]. At the distributional pole, the probability density function resembles a Delta distribution.
@@
In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ more when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero. This resemblance allows for a generic characterization of the output.
@@ 2.3 Statistical Taylor Expansion
Formula (2.2) provides the uncertainty distribution of an analytic function. However, in most scientific and engineering calculations, the primary interest is not the full result distribution, but rather just few summary statistics of the result, such as the mean and deviation [2][3]. These simplified statistics can be obtained through a statistical Taylor expansion.
Let $\overline{x}$ and $\delta x$ denote the sample mean and deviation of a random variable. Let $\mu$ and $\sigma$ denote the distribution mean and deviation of the same variable. An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.5). Formula (2.4) defines the bound moment $\zeta(n,\kappa)$, where $0<\varrho,\kappa$ specify the bounding ranges. Using Formula (2.4), Formula (2.6) and Formula (2.3) yields the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the result value.
@@
Formula (2.4) defines a bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ for the underlying distribution, where $\mu$ and $\sigma$ are the distribution mean and distribution deviation, respectively. As discussed later in this paper, $\kappa$ statistically determines $\varrho$, while $\kappa$ itself is determined by both the sample size and the underlying distribution of the input. Certain $f(x)$ such as $\log(x)$ and $x^{c}$ also require an upper bound on $\kappa$ to ensure the convergence of Formula (2.6) and (2.3). The probability $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
Under the uncorrelated uncertainty assumption, Formula (2.9) and (2.3) compute the mean and variance of the Taylor expansion in Formula (2.8), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ are the variance moments for $x$ and $y$, respectively:
@@
Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
@@ 2.4 Bounding Symmetry
A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbound distribution. To achieve mean preserving, $\kappa$ determines $\varrho$, so that $\zeta(n)$ becomes $\zeta(n,\kappa)$. Under a mean-preserving bounding, Formula (2.11) and (2.12) provide the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
@@
When $4\leq\kappa$, $\zeta(0,\kappa)\simeq 1$ and $\zeta(2,\kappa)\simeq 1$, making Formula (2.11) and (2.12) the convolution results for $x\pm y$ [5], and Formula (2.13) and (2.14) the corresponding results of the product distribution for $xy$ [5].
For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
@@ 2.5.1 Uniform Input Uncertainty
$\eta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.15)
@@ 2.5.2 Gaussian Input Uncertainty
The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [5]. This convergence occurs rapidly [1]. In a digital computer, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sum of expansion terms [8][9]. Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [2][3][5].
For Gaussian input uncertainty, $\zeta(2n+1)=0$.
If $\kappa=\infty$, then $\zeta(2n)=(2n-1)!!$, which causes Formula (2.29) for $\log(x\pm\delta x)$ and Formula (2.35) for $(x\pm\delta x)^{c}$ to diverge. As discussed, $\kappa$ is determine by the sample count $N$. When $\kappa$ has an upper bound, Formula (2.4) reduces to Formula (2.16):
@@
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$ according to Formula (2.18). When $\kappa=5$, and $n<5$, the relative error satisfies $\left|\eta(2n)/(2n-1)!!-1\right|<10^{-3}$.
@@ 2.5.3 An Input Uncertainty with Limited Range
$\zeta(n)$ $=(n+1)!\left(e^{-\varrho}\sum_{j=0}^{n+1}\frac{\varrho^{j}}{j!}-e^{-\kappa}\sum_{j=0}^{n+1}\frac{\kappa^{j}}{j!}\right);$ (2.22)
$\lim_{n\rightarrow+\infty}\zeta(n)=\frac{\kappa^{n+2}}{n+2}e^{-\kappa};$ (2.23)
Formula (2.20) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<1<\kappa$. Formula (2.21) gives its mean-preserving equation. Formula (2.22) expresses its bound moment. And Formula (2.23) describes its asymptotic behavior.
@@ 2.8 Convergence
Formula (2.26) for $e^{x\pm\delta x}$ and Formula (2.32) for $\sin(x\pm\delta x)$ both converge unconditionally. However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can take negative values for large $\delta x$, which imposes a constraint upper bound on the input $\delta x$.
Formula (2.29) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) when $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$ for a Gaussian distribution, or $P(x)<1/\sqrt{3}$ for a uniform distribution.
@@
Formula (2.35) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.44) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ . This expression converges when $P(x)\lesssim 1/\kappa$ although the precisse upper bound for $P(x)$ varies with $c$.
@@
2.9 Bounding Range
Figure 3: The result variance ratio $\delta^{2}f/\widehat{\delta^{2}f}$ (left y-axis) and the bounding leakage (right y-axis) for varying bounding range $\kappa$ (x-axis) for the selected function $f(x)$ (legend), where $\widehat{\delta^{2}f}$ denotes the corresponding variance for $\kappa=5$.
Figure 4: The ratio of adjusted $\widehat{\delta^{2}f}$ to the original $\widehat{\delta^{2}f}$ (as shown by the y-axis) for varying bounding range $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), where $\widehat{\delta^{2}f}$ is approximated by normalizing the input variance with $\zeta(2,\kappa)$ such that $\delta^{2}x=(\delta x)^{2}$.
The bounding range $\kappa$ influences the results of statistical Taylor expansion unless it is sufficiently large. Figure 3 shows that $\epsilon(\kappa)$ decreases as $\kappa$ increases, while $\delta^{2}f$ for the selected functions rises with $\kappa$ until reaching stable value when $\kappa\geq 4$. When $\kappa\geq 5$, the results show negligible difference for varying $\kappa$. In practice, based on the 5-$\sigma$ rule [3] for reliably distinguishing two random variables, let $\widehat{\delta^{2}f}$ denote the ideal variance obtained once the result variance stabilizes for $\kappa=5$. Figure 3 indicates that $\delta^{2}f\leq\widehat{\delta^{2}f}$ due to $\epsilon(\kappa)$, and when $\epsilon(\kappa)\simeq 0$, $\delta^{2}f\simeq\widehat{\delta^{2}f}$. The variance ratio is defined as $\delta^{2}f/\widehat{\delta^{2}f}$. Figure 3 also demonstrates that variance ratio remains approximately a constant for a $\kappa$ value.
To approach $\widehat{\delta^{2}f}$ when $\kappa<5$, instead of $(\delta x)^{2}$, the adjusted input uncertainty variance $(\delta x)^{2}/\zeta(2,\kappa)$ is used in Formula (2.3) to obtain the adjusted variance whose $(\delta x)^{2}$-term has the value as if $\zeta(2,\kappa)=1$, which is same as that of the ideal variance $\widehat{\delta^{2}f}$. Figure 4 shows that the adjusted variances are close to the corresponding ideal variance, however as a first order approximation, their ratios depend on the specific form of $f(x)$. Figure 4 also indicates a universal worst $\kappa\simeq 1.75$, below which the tail effect of Gaussian distributions vanishes.
$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.45)
$\zeta(2,\kappa)\equiv$ $\frac{\delta^{2}f}{\widehat{\delta^{2}f}}$; (2.46)
$x\pm y:$ $\;\;\;\zeta(2,\kappa)=\frac{\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.47)
Formula (2.45) is a special case for Formula (2.3), which can be reinterpreted as obtaining the bounding range $\kappa$ from a variance ratio. This relation between variance ratio and bounding range can be generalized as in Formula (2.46). For example, Formula (2.47) shows that the result $\kappa$ for $x\pm y$ is a variance-weighted average of the input $\kappa_{x}$ and $\kappa_{y}$ via $\zeta(2,\kappa)$.
@@ 2.10 Statistical Bounding
Figure 6: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range for the normal distribution.
@@
[…]
When the underlying distribution is uniform, $\kappa_{s}=\sqrt{3}$ and $\epsilon=0$. Figure 6 shows that $0<\epsilon(N)\sim N^{-0.564}$, resulting in a measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.4).
In Figure 6, when the underlying distribution is Gaussian, the bounding range $\kappa(5,N)$ start to reach it stable value at sample count $N=10^{3}$. In contrast, when the underlying distribution is uniform, such sample count $N$ is $10^{2}$. This difference of the sample count upper bounds necessary for computing adjusted variances is caused by the underlying distributions.

### v12

@@ 2.2 Distributional Zero and Distributional Pole
$\rho(\tilde{x},\mu,\sigma)d\tilde{x}$ $=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y}$; (2.5)
@@
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, since $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [5]. A linear transformation generates neither a distributional zero nor a distributional pole, according to Formula (2.5).
  • Figure 3 shows the probability density function for $(x\pm 1)^{2}$ according to Formula (2.6), which exhibits a distributional pole at $x=0$. The distribution $(0\pm 1)^{2}$ corresponds to the $\chi^{2}$ distribution [2]. At the distributional pole, the probability density function resembles a Delta distribution.
@@
In both Figure 3 and 4, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, thereby allowing for a generic characterization of the output.
@@ 2.3 Statistical Taylor Expansion
Formula (2.5) provides the uncertainty distribution of an analytic function. However, in most scientific and engineering calculations, the primary interest lies not in the full result distribution but in few summary statistics of the result, such as the mean and deviation [2][3]. These simplified statistics can be obtained through a statistical Taylor expansion.
[…]
@@
An confidence interval [5] $\tilde{x}\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ can describe a sampling from an underlying distribution, where $0<\varrho,\kappa$ specify the bounding ranges. Formula (2.7) defines the bound moment $\zeta(n)$. As discussed later, $\kappa$ determines $\varrho$, while $\kappa$ itself is determined by both the sample size $N$ and the underlying distribution of the input. When $N\rightarrow\infty$, $\varrho,\kappa\rightarrow\infty$, so that by definition $\zeta(0)=1$, $\zeta(1)=0$, and $\zeta(2)=1$. When $N$ is bounded, $\varrho$ and $\kappa$ are also bounded, so that $\zeta(0)<1$, and $\zeta(2)<1$. The probability of $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.8). Using Formula (2.7), Formula (2.9) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
@@
Under the uncorrelated uncertainty assumption, Formula (2.12) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.11), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively. Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
@@ 2.4 Bounding Symmetry
A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbounded distribution. To achieve mean preserving, $\kappa$ determines $\varrho$, so that $\zeta(n)$ becomes $\zeta(n,\kappa)$. Under a mean preserving bounding, Formula (2.14) and (2.15) provide the result for $x\pm y$, while Formula (2.16) and (2.17) give the result for $xy$:
@@
When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.14) and (2.15) the convolution results for $x\pm y$ [5], and Formula (2.16) and (2.17) the corresponding results of the product distribution for $xy$ [5].
For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion. Both Gaussian and Uniform distributions are symmetric.
@@ 2.5.1 Uniform Input Uncertainty
$\zeta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.18)
@@ 2.5.2 Gaussian Input Uncertainty
The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [5]. This convergence occurs rapidly [1]. In digital computation, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sums of expansion terms [8][9]. Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [2][3][5].
Formula (2.7) reduces to Formula (2.19) and (2.20):
@@
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$ according to Formula (2.21). When $\kappa=5$, and $n<5$, the relative error $\left|\zeta(2n)/(2n-1)!!-1\right|$ is less than $10^{-3}$.
@@ 2.5.3 An Input Uncertainty with Limited Range
$\lim_{n\rightarrow+\infty}\zeta(n,\kappa)$ $=\frac{\frac{\kappa}{\sqrt{2}}^{n}}{n+2}e^{-\kappa};$ (2.25)
Formula (2.23) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<2<\kappa$. Formula (2.24) gives its mean preserving equation. And Formula (2.25) describes its asymptotic behavior.
@@ 2.8 Convergence
Formula (2.28) for $e^{x\pm\delta x}$ and Formula (2.34) for $\sin(x\pm\delta x)$ both converge unconditionally. However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can become negative for large $\delta x$, which imposes an upper-bound constraint on the input $\delta x$.
Formula (2.31) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) as $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$.
@@
Formula (2.37) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.46) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ . This expression converges when $P(x)\lesssim 1/\kappa$ although the precis upper bound for $P(x)$ varies with $c$.
@@ 2.9 Statistical Bounding
Figure 6: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range $\kappa_{s}$ for the normal distribution.
@@
When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to bounding leakage $\epsilon(N)$. Figure 6 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically. The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.7).
@@
When the underlying distribution is Normal, Formula (2.47) and (2.48) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Gaussian error function [5]. Figure 5 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$. Figure 6 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$). Figure 6 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$. Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 5 should be used as $\kappa$ in Formula (2.7).
2.10 Ideal Statistics
Figure 7: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Gaussian.
Figure 8: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Uniform.
$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.49)
$\alpha(x\pm y)=$ $\frac{\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.50)
When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}\rightarrow\sigma$ in Formula (2.49). Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
  • When the underlying uncertainty distribution is Gaussian, Figure 7 shows that the resulting variance $\delta^{2}f$ for a selected functions rise with NN until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 20$. When $N=20$ and $\kappa_{s}=5$, $\kappa=3.8$ according to Figure 5. Therefore it is sufficient to compute the stable variances $\widehat{\delta^{2}}f$ with $\zeta(n,5)$.
  • When the underlying uncertainty distribution is Uniform, Figure 8 shows the same trend, however the stable variances are reached only when $N>10^{5}$. The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.18).
Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input. Ideal variances are consistent with traditional variances [5].
In both Figure 7 and 8, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding ideal variances with exactly the same trend. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$. Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution. When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.50) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
An ideal ratio also applies to the corresponding uncertainty bias. The trends of uncertainty biases $\overline{f}-f$ versus sample count NN are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.